OpenStudy (anonymous):

Solve by substitution: x + y = 3 2x + 3y = -4

OpenStudy (anonymous):

more info needed here? what I am to sub in order to solve?

OpenStudy (anonymous):

x + y = 3...(i) 2x + 3y = -4....(ii) from (i) x=3-y substitute in (ii) 2(3-y) +3y = -4 6-2y+3y=-4 y=-4-6 y=?? x=??

OpenStudy (anonymous):

so does x = -10 ?

OpenStudy (anonymous):

no that is y = -10

OpenStudy (anonymous):

so how do i get x ?

OpenStudy (anonymous):

-10 and sub in for x and slove

OpenStudy (anonymous):

@breeskee03 y=-10 remember from equation (i) x=3-y so just substitute to get x

OpenStudy (anonymous):

so x= -7

OpenStudy (anonymous):

@godorovg that's incorrect x=3-y x=3--10 x=3+10 x=13

OpenStudy (anonymous):

opps dropped a minus sign sorry yes x=13